(1/2 x+4) + (3/4 x-6)=2 x-47
Found 2 solutions by solver91311, macston:
Answer by solver91311(24713)



What does (1/2 x+4) mean?

or

What does (3/4 x-6) mean?

or

What does 2 x-47 mean?

or

John

My calculator said it, I believe it, that settles it

Answer by macston(5194)

(1/2 x+4) + (3/4 x-6)=2 x-47
(2/4)x+(3/4)x+4-6=2x-47
(5/4)x-2=2x-47
45=(3/4)x
60=x
.
CHECK:
.
(1/2 x+4) + (3/4 x-6)=2 x-47
(1/2(60)+4)+(3/4(60)-6)=2(60)-47
(30+4)+(45-6)=120-47
34+39=73
73=73